OpenStudy (anonymous):

can someone kindly tell me if we can split this : (3-6t)+1 / (3-6t)(3-6t) into....... (3-6t)/(3-6t) + 1/(3-6t) ??

sam (.sam.):

\[\frac{(3-6t)+1}{(3-6t)(3-6t) }\] \[\frac{(3-6t)}{(3-6t)(3-6t)} +\frac{1}{(3-6t)(3-6t)}\] \[\frac{1}{(3-6t)}+\frac{1}{(3-6t)(3-6t)}\]

sam (.sam.):

Unless from\[\frac{1}{(3-6t)}+ \frac{1}{(3-6t)(3-6t)}\] divide whole thing by (3-6t), then \[\frac{(3-6t)}{(3-6t)} +\frac{1}{(3-6t)}\]

OpenStudy (anonymous):

that means the answer is yes

sam (.sam.):

yes but normally for \[\frac{(3-6t)}{(3-6t)}\] is written as "1"